Hi, please help me solve this using the systems of equations. Thank you!
The perimeter of a rectangle rug is 40 ft and the area is 96 ft^2. Find the dimensions of the rug.

Answer by ikleyn(52776)

.
The system of equations is 

2L + 2W = 40,
LW      = 96.

Or

L + W = 20,
LW    = 96.


Answer. 12 and 8.

Solve by substitution
L = 20- W,

(20-W)*W = 96, or

W^2 - 20W + 96 = 0,

(W-12)*(W-8) = 0.

Etc. . . .